# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ II.2 Algorithm
Simulating long-range interactions using the Monte Carlo method presents two key challenges: the critical slowing-down near phase transitions and escalating simulation time with increasing system size. Critical slowing-down arises near the critical point, where consecutive samples become highly correlated. This correlation leads to a significant decrease in simulation efficiency as the system size increases. Various strategies have been developed to address this issue, such as cluster algorithms [27, 28], direct-loop algorithms [29], and worm algorithms [30]. The rapidly increasing simulation time is another major hurdle in accurately analyzing the critical properties of a long-range system. For a long-range interacting spin system of size $N$, each spin interacts with other $N-1$ spins, so a proposed local MC update requires evaluating all $N-1$ interaction terms. This results in a computational complexity $\mathcal{C}$ of $\mathcal{O}(N)$ for the LR system, compared to $\mathcal{O}(1)$ for the NN case. Here, the computational complexity is defined as the average number of operations required to update a single spin, serving as a measure of the time complexity of the MC algorithm. In practice, the actual CPU time of an MC local update is proportional to $\mathcal{C}$. Several methods have been proposed to mitigate this challenge. Techniques such as the worm algorithm combined with the diagrammatic Monte Carlo method expand the pairwise potential energy into diagrammatic contributions, making the computational complexity independent of system size [31]. Efficient cluster algorithms for the classical long-range Ising model have also been developed [32, 33]. The recently introduced clock Monte Carlo method [26, 34], based on the factorized Metropolis filter [35], reduces computational complexities of several classical LR models to $\mathcal{O}(1)$.
@@
In this work, we simulate the LR XY model using an enhanced variant of the Luijten-Blöte (LB) algorithm [32, 36]. We augmented the original LB algorithm with the clock Monte Carlo method and extended it to the $O(N)$ spin model. This approach employs the cluster updates scheme to address the issue of critical slowing down and adopts an efficient cluster construction strategy to reduce the computational complexity associated with long-range interactions. Without loss of generality, we elucidate the algorithm for the $O(\mathcal{N})$ spin model with long-range interactions.
@@
We begin with the standard multi-cluster algorithm for a long-range $O(\mathcal{N})$ spin model. A Monte Carlo sweep consists of two main steps: cluster construction and cluster flipping. The cluster construction starts by randomly choosing a unit vector $\mathbf{S}_{\rm{ref}}$ in the $\mathcal{N}$-dimensional space. Each spin $\mathbf{S}$ is projected onto this reference vector, yielding the parallel component $S^{\parallel}=\mathbf{S}\cdot\mathbf{S}_{\rm{ref}}$. Spin clusters are then formed by sequentially inspecting all pairs of spins (bonds). Each bond $(i,j)$ is independently activated with probability,
@@
with $[x]^{+}={\rm max}(0,x)$. After all the bonds have been examined, spins connected by active bonds form clusters, which are then independently flipped with probability $1/2$. A flipped spin is given by $\bm{S}^{\prime}=\bm{S}-2S^{\parallel}\bm{S}_{\rm{ref}}$. An updated configuration is generated after all clusters have been considered for flipping. The algorithm operates with a computational complexity of $\mathcal{O}(N)$, as each MC sweep involves evaluating $N(N-1)/2$ bonds, and on average, $\mathcal{O}(N)$ operations are required per spin.
In the 1990s, Luijten and Blöte introduced an efficient cluster algorithm for the long-range Ising model [32]. Their method is based on the observation that, on average, only $O(N)$ among $O(N^{2})$ bonds participate in the cluster construction. They proposed a rejection-free approach using the binary search to directly sample the next bond to activate (the bond activation event) instead of sequentially inspecting all bonds. This strategy reduces the number of operations per MC sweep to $O(N\log N)$, resulting in a computational complexity of $O(\log N)$, which drastically increases the efficiency of the algorithm. Nevertheless, their method is primarily limited to the LR Ising model. Additionally, the binary search requires a look-up table to sample the bond activation events, making it technically challenging to apply to 2D models, and special modifications to the coupling function are needed [8].
To overcome these limitations, we integrated the clock Monte Carlo method with the LB algorithm. The clock Monte Carlo method replaces the conventional Metropolis acceptance criterion with a factorized Metropolis filter — a product of independent factors where an update is rejected if any independent factor fails. By directly sampling the first rejecting factor (first-rejection event) using the clock sampling technique, the computational complexity is significantly reduced from $O(N)$ to $O(1)$. Recognizing the similarity between first-rejection events and bond activation events, we apply the clock sampling technique to sample bond activation events efficiently.
@@
This probability depends solely on the strength of the interaction $J_{ij}$ and ensures that for any spin configuration, $\hat{\rho}_{ij}\geq\rho_{ij}$. Hence, $\hat{\rho}_{ij}$ is referred to as the bound activation probability. The actual activation of a bond $(i,j)$ is then determined by first attempting a bound activation with probability $\hat{\rho}_{ij}$, followed by resampling with a relative probability:
@@
Therefore, by introducing the bound activation probability and the resampling process, a bond inspection is only required when a bound acceptance occurs. We shall leverage this property for efficient sampling of bond activations.
In a long-range $O(\mathcal{N})$ model, the bonds can be categorized into different types based on their displacement vector $\vec{r}$ [37]. Each type of bond shares the same bound activation probability $\hat{\rho}$. To construct clusters, we first obtain, for each type of bond, a list of bonds that are bound activated. Then, these bound-activated bonds are resampled to inspect for actual activation. Such a list can be obtained by sampling the distribution $P(k)=\hat{\rho}(1-\hat{\rho})^{k-1}$ which expresses the probability that the $k$-th bonds are bound activated, while $k-1$ previous bonds are rejected and skipped. When the $\hat{\rho}$ is small, which is the case near $T_{c}$, most bonds are skipped before a bond is inspected for activation. For a type of bond that consists of $N_{k}$ bonds with $\hat{\rho}$, the bond activation process works as follows:
@@
$k\leftarrow k+1+\left\lfloor\frac{\ln(r)}{\ln(1-\hat{\rho})}\right\rfloor,$ (9)
where $r$ is a uniform random number in $(0,1]$ and $\lfloor x\rfloor$ denotes the floor function that returns the integer part of $x$.
(c) If $k\leq N_{k}$, attempt to activate the $k$-th bond with relative probability $\rho_{k,\text{rel}}$.
(d) Repeat from step (b) until $k>N_{k}$.
The cluster construction process concludes once all types of bonds have been considered. This procedure leads to a considerably faster algorithm due to the reduced number of bond inspections and random number generation. Compared to the LB algorithm, our method eliminates the need for a look-up table and alleviates truncation errors associated with discrete cumulative probability integration approximations. It reproduces the stochastic dynamics of the original multi-cluster method while achieving a $O(1)$ computational complexity within the parameter space studied in this work. In practice, care should be taken to avoid double-counting, which can be achieved by selecting unique displacement vectors on the torus.
We have performed extensive MC simulations of the LR XY model using the algorithm described in this subsection, with system sizes ranging from $L=16$ to $L=8192$. Each system is thermalized for $4\times 10^{3}$ MC sweeps, which is sufficient to reach thermal equilibrium. For system sizes $L\leq 4096$ at low temperatures $\beta=1,2,4,8$ and near the critical temperature of various $\sigma$, we collect over $2\times 10^{6}$ samples to achieve high statistical accuracy. Under other parameters, where high-precise data is not required for fitting, the minimal number of samples is $4\times 10^{5}$. Simulations for the largest system size $L=8196$ are conducted at selected parameters, and the results are obtained from over $4\times 10^{4}$ samples. Like other cluster algorithms, our simulation does not suffer from the critical-slowing down issue. Specifically, for a simulation with $L=8192$ and $\sigma=2$ near the critical point $T_{c}$, the integrated autocorrelation time $\tau_{\text{int}}$ in the unit of MC sweep is $\sim 9$ for the energy-like quantity and $\sim 4.6$ for the squared magnetization. In this particular case, the average computational complexity is $\sim 2$ operations per spin, which remains virtually constant for all simulations near $T_{c}$ in this work, and it takes around 11 CPU days to go over $10^{4}$ MC sweeps. Moreover, the CPU time per spin update is proportional to the computational complexity, and the space complexity of the algorithm scales linearly with system volume $N$. For the largest system size $L=8192$ simulated in this work, our implementation requires $\sim 10.5$ Gigabyte of memory.
@@ II.3 Sampled Quantities
    Magnetization density $\mathbf{M}=L^{-2}\sum_{i=0}^{N}\bm{S}_{i}$, and its Fourier mode $M_{k}=L^{-2}\left|\sum_{i}\bm{S}_{i}e^{i\bm{k}\cdot\bm{r}_{i}}\right|$, where $\bm{k}=\frac{2\pi}{L}\hat{\bm{x}}$, the smallest wave vector of the reciprocal lattice in $\bm{x}$-direction.
@@
    $Q_{m}=\frac{\langle M^{2}\rangle^{2}}{\langle M^{4}\rangle}$ (11)
@@
    The scaled covariance between $\varepsilon$ and $M^{2}$,
    $K$ $=-\frac{L^{2}}{\langle M^{2}\rangle}\left(\langle\varepsilon M^{2}\rangle-\langle\varepsilon\rangle\langle M^{2}\rangle\right)$ (12)
Note that the second-moment correlation length $\xi_{\text{2nd}}$ is different from the commonly referred exponential correlation length $\xi_{\text{exp}}$. The former is defined based on the Fourier spectrum of the correlation function [39, 41], while the latter is obtained by fitting the exponentially decaying correlation function. Both correlation lengths are well-defined in the disordered phase and become asymptotically equivalent in the thermodynamic limit, sharing the same scaling behavior near the critical point. Compared to $\xi_{\text{exp}}$, the second-moment correlation $\xi_{\text{2nd}}$ is much easier to measure in MC simulations. Conversely, obtaining $\xi_{\text{exp}}$ can be challenging due to finite-size effects that complicate the fitting process. Hereafter, unless specified otherwise, the correlation length $\xi$ refers to the second-moment correlation length.
Furthermore, the dimensionless ratio $\xi/L$ is an effective tool for identifying the critical points of phase transitions [42, 43, 40]. At criticality, quantity $\xi/L$ attains a universal value in the thermodynamic limit. In the disordered phase, where the correlation length $\xi$ is finite, this ratio drops to zero as the system size $L$ increases. In contrast, in the ordered phase, $\xi/L$ diverges rapidly because $M^{2}$ remains finite while $M_{k}^{2}$ vanishes quickly with increasing $L$. For a second-order transition, the intersection of $\xi/L$ curves for different $L$ accurately pinpoints the critical point. In the case of a BKT transition, however, the ratio converges to a universal curve in the QLRO phase.
Throughout the simulation, the error bars are estimated using standard error analysis techniques: for simple observables, such as $\varepsilon,M^{2}$, and $M_{k}^{2}$, the error bars are calculated using the standard binning method; for composite observables obtained by the arithmetic combinations of simple observables, such as the 2nd-moment correlation length $\xi$ and Binder ratio $Q$, the error bars are estimated using the jackknife method.
@@ II.4 Finite-size Scaling Analysis
Finite-size scaling (FSS) analysis is an essential technique in the numerical investigation of critical phenomena. In the language of the renormalization group (RG), the critical behavior of systems belonging to the same universality class is governed by a common fixed point. Near critical points, the free energy, and thus other physical observables, can be expressed as universal functions of scaling fields. FSS extends these universal functions to finite systems, enabling precise determination of critical points and accurate extraction of critical exponents. Near the critical point, the free-energy density $f(t,h,u,L^{-1})$ is a function of the thermal scaling field $t$, the magnetic scaling field $h$, an irrelevant field $u$, and the finite-size field $L^{-1}$. Then, the FSS behavior of free-energy density is given by,
@@
where $y_{t}$, $y_{h}$ and $y_{i}$ are the corresponding scaling exponent and $g$ is the analytic part of free-energy density. The scaling fields are non-universal functions of microscopic interactions and external physical parameters. To the lowest order, $t$ is proportional to $T-T_{c}$, i.e., the distance between temperature and critical temperature; $h$ is proportional to the physical external magnetic field $H$. The irrelevant field $u$ reflects the distance of criticality of the system and the fixed point.
The FSS behavior of various physical observables near $T_{c}$ can be derived from Eq. (13). Here, we present the scaling form of several key quantities near or at a second-order phase transition: (a) magnetic susceptibility $\chi$, (b) scaled covariance $K$, and (c) dimensionless quantities, such as $Q_{m}$ and $\xi/L$.
(a) Scaling behavior of $\chi$. The magnetic susceptibility $\chi$ is calculated from the averaged squared magnetization $\langle M^{2}\rangle$, which is propositional to $\partial^{2}f/\partial h^{2}$ at $h=0$. Thus, near the critical point, the magnetic susceptibility scales as
@@
(b) Scaling behavior of $K$. The quantity $K$ is proportional to the correlation between the squared magnetization $M^{2}$ and the NN energy $\varepsilon$. The NN energy $\varepsilon$ shares a similar scaling behavior as the total energy $E$ near $T_{c}$ as $\sim L^{y_{t}-d}$ [37, 44]. Consequently, the correlation between $M^{2}$ and $\varepsilon$ is proportional to the temperature derivative of squared magnetization $\frac{\partial}{\partial t}\langle M^{2}\rangle$ [37]. The FSS form of $K$ at the critical point $t=0$ is then given by
@@
where $a_{i}$ are non-universal constants. The leading $L^{y_{t}}$ scaling behavior in Eq. (16) makes $K$ an excellent quantity to estimate the thermal scaling exponent $y_{t}$.
(c) Scaling behavior of dimensionless ratios. Dimensionless ratios, such as Binder ratio $Q_{m}$ and correlation length ratio $\xi/L$, have been widely used in FSS analysis to locate the critical point of phase transitions [42, 45, 43]. Near the critical point, the Binder ratio follows the scaling behavior
@@
III Results: Overview
Figure 1: The phase diagram of the two-dimensional long-range XY model reveals distinct behaviors depending on the interaction range $\sigma$ and the inverse temperature $\beta=1/T$. In the short-range regime ($\sigma>2$), the system exhibits BKT transitions (brown line) into a QLRO phase. In the non-classical regime ($1<\sigma\leq 2$), the system undergoes a continuous phase transition (red line) into a long-range ordered (LRO) phase. Lastly, in the classical regime ($\sigma\leq 1$), the transition (purple lines) is governed by Gaussian theory. The symbols $\beta_{c}^{\text{CG}}$ and $\beta_{c}^{\text{NN}}$ denote the critical inverse temperatures for the complete-graph and NN interaction limits, respectively.

### v2

@@ II.2 Algorithm
Simulating long-range interactions using the Monte Carlo method presents two major challenges: critical slowing down near phase transitions and escalating computational costs with increasing system sizes. Critical slowing down occurs near the critical point, where successive samples become highly correlated, resulting in a substantial reduction in simulation efficiency as system size increases. Various update schemes have been developed to address this issue, such as cluster algorithms [28, 29], direct-loop algorithms [30], and worm algorithms [31]. The second challenge stems from the nature of long-range interactions: each spin interacts with the other $N-1$ spins, so a proposed local MC update requires evaluating all $N-1$ interaction terms. This results in a computational complexity $\mathcal{C}$ of $\mathcal{O}(N)$ for the LR system, in contrast to $\mathcal{O}(1)$ for the NN case. We define the computational complexity as the average number of operations required to update a single spin, serving as a measure of the time complexity of the MC algorithm. In practice, the CPU time per local update scales linearly with $\mathcal{C}$. Several methods have been proposed to mitigate this cost. For example, the worm algorithm combined with the diagrammatic Monte Carlo method expands the pairwise potential energy into diagrammatic contributions, making the computational complexity independent of the system size [32]. Efficient cluster algorithms tailored for the classical long-range Ising model have also been developed [33, 34]. More recently, the clock Monte Carlo method [35, 36], based on the factorized Metropolis filter [37], reduces computational complexities of several classical LR models to $\mathcal{O}(1)$, thus offering a scalable and efficient solution to simulating large-scale LR systems.
@@
We simulate the LR XY model using a variant of the Luijten-Blöte (LB) algorithm [33, 38]. We enhance the original LB algorithm with the clock Monte Carlo method and extend it to the $O(n)$ spin model. This approach employs the cluster updates scheme to address the issue of critical slowing down and adopts an efficient cluster construction strategy to reduce the computational complexity associated with long-range interactions.
@@
We begin with the standard Swendson-Wang algorithm for a long-range $O(n)$ spin model [28]. A Monte Carlo sweep consists of two main steps: cluster construction and spin flipping. The cluster construction starts by randomly choosing a unit vector $\mathbf{S}_{\rm{ref}}$ in the $n$-dimensional space. Each spin $\mathbf{S}$ is projected onto this reference vector, yielding the parallel component $S^{\parallel}=\mathbf{S}\cdot\mathbf{S}_{\rm{ref}}$. Only this projection component can undergo spin flipping during the update process, while the remaining orthogonal components remain invariant. This is effectively an Ising-type update scheme with the coupling constant between $\boldsymbol{S}_{i}$,$\boldsymbol{S}_{j}$ as $J_{ij}S^{\parallel}_{i}S_{j}^{\parallel}$. Spin clusters are then formed by sequentially inspecting all pairs of spins (bonds). Each bond $(i,j)$ is independently activated with probability,
@@
with $[x]^{+}={\rm max}(0,x)$ and the inverse temperature $\beta=1/T$. After all the bonds have been examined, spins connected by active bonds form clusters, which are then independently flipped with probability $1/2$. A flipped spin is given by $\bm{S}^{\prime}=\bm{S}-2S^{\parallel}\bm{S}_{\rm{ref}}$. An updated configuration is generated after all clusters have been considered. The algorithm operates with a computational complexity of $\mathcal{O}(N)$, because each MC sweep involves evaluating $N(N-1)/2$ bonds, and on average, $\mathcal{O}(N)$ operations are required per spin.
In the 1990s, Luijten and Blöte introduced an efficient cluster algorithm for the long-range Ising model [33]. Their method is based on the observation that, on average, only $\mathcal{O}(N)$ among $\mathcal{O}(N^{2})$ bonds participate in the cluster construction. They proposed a rejection-free approach using the binary search to directly sample the next bond to activate (the bond activation event) instead of sequentially inspecting all bonds. This strategy reduces the number of operations per MC sweep to $\mathcal{O}(N\log N)$, resulting in a computational complexity of $\mathcal{O}(\log N)$, which drastically increases the efficiency of the algorithm. Nevertheless, their method is primarily limited to the LR Ising model. Additionally, the binary search requires a look-up table to sample the bond activation events, making it technically challenging to apply to 2D models, and special modifications to the coupling function are needed [6].
To overcome these limitations, we integrate the clock Monte Carlo method with the LB algorithm. The clock Monte Carlo method replaces the conventional Metropolis acceptance criterion with a factorized Metropolis filter, which expresses the acceptance probability as a product of independent factors; an update is rejected if any individual factor fails. By directly sampling the first rejecting factor (first-rejection event) using the clock sampling technique, the computational complexity is significantly reduced from $\mathcal{O}(N)$ to $\mathcal{O}(1)$ [35]. Recognizing the similarity between first-rejection events and bond activation events, we apply the clock sampling technique to efficiently sample bond activations.
@@
This probability depends solely on the strength of the interaction $J_{ij}$ and ensures that for any spin configuration, $\hat{\rho}_{ij}\geq\rho_{ij}$. Accordingly, $\hat{\rho}_{ij}$ is referred to as the bound activation probability and activating a bond according to $\hat{\rho}_{ij}$ is called a bound activation. The actual activation of a bond $(i,j)$ then proceeds in two steps: a bound activation is first attempted with probability $\hat{\rho}_{ij}$; if successful, one performs a resampling with the relative probability:
@@
Through this two-step process, an explicit evaluation of the bond condition Eq. (4) is only required when a bound activation occurs. We exploit this property to achieve efficient sampling of bond activations.
In a long-range $O(n)$ spin model, there are $(N-1)/2$ types of displacement vectors $\vec{r}$. For each $\vec{r}$, there are $N$ interactions $J(\vec{r})$ due to the translational invariance of the system with PBC, composing a total of $N(N-1)/2$ interaction pairs [39]. All bonds corresponding to the same $\vec{r}$ then share an identical $\hat{\rho}$. To construct clusters, we first generate, for each type of $\vec{r}$, a list of bound-activated bonds. Then, these bonds are resampled to determine the actual activation. The list can be efficiently generated by sampling the distribution $P(k)=\hat{\rho}(1-\hat{\rho})^{k-1}$, which is the probability that the $k$-th bonds are bound activated, while the preceding $k-1$ bonds are rejected and skipped. This method is particularly advantageous when the $\hat{\rho}$ is small, which is typically the case near $T_{c}$; most bonds are skipped before a bond is inspected for activation, significantly reducing the overall computational cost. For a type of bond that consists of $N$ bonds with $\hat{\rho}$, the bond activation process works as follows:
@@
$k\leftarrow k+1+\left\lfloor\frac{\ln(\texttt{rand})}{\ln(1-\hat{\rho})}\right\rfloor,$ (7)
where rand is a uniform random number in $(0,1]$ and $\lfloor x\rfloor$ denotes the floor function that returns the integer part of $x$.
(c) If $k\leq N$, attempt to activate the $k$-th bond with relative probability $\rho_{k,\text{rel}}$ in Eq, (6).
(d) Repeat from step (b) until $k>N$.
The cluster construction process concludes once all types of bonds have been considered. This procedure leads to a considerably faster algorithm by reducing both the number of bond inspections and the number of random number generations. Compared to the LB algorithm, our method eliminates the need for a look-up table and alleviates truncation errors associated with discrete cumulative probability integration approximations. It faithfully reproduces the stochastic dynamics of the original multi-cluster method while achieving $\mathcal{O}(1)$ computational complexity.
In this work, we have performed extensive MC simulations of the LR XY model using the enhanced LB algorithm, with linear system sizes ranging from $L=16$ to $L=8192$. Each system is thermalized for $4\times 10^{3}$ MC sweeps, which is sufficient to reach thermal equilibrium. For system sizes $L\leq 4096$ at low temperatures $\beta=1,2,4,8$ and near the critical temperature of various $\sigma$, we collect over $2\times 10^{6}$ samples to achieve high statistical accuracy. Under other parameters, where high-precise data is not required, the minimal number of samples is $4\times 10^{5}$. Simulations for the largest system size $L=8196$ are conducted at selected parameters, and the results are obtained from over $4\times 10^{4}$ samples. Our simulation for the 2D LR XY model near $\sigma=2$ does not encounter strong critical slowing down issues. Specifically, for a simulation with $L=8192$ and $\sigma=2$ near the critical point $T_{c}$, the integrated autocorrelation time $\tau_{\text{int}}$ in the unit of MC sweep is $\sim 9$ for the energy-like quantity and $\sim 4.6$ for the squared magnetization. The average computational complexity is $\sim 2$ operations per spin, which remains virtually constant for all simulations near $T_{c}$, and it takes around $11$ CPU days to go over $10^{4}$ MC sweeps for a CPU with frequency $2.4$GHz. Moreover, the CPU time per spin update is proportional to the computational complexity, and the space complexity of the algorithm scales linearly with system volume $N$. For the largest system size $L=8192$, our implementation requires $\sim 10.5$ gigabytes of memory.
@@ II.3 Sampled Quantities
    Magnetization density $\mathbf{M}=L^{-2}\sum_{i=0}^{N}\bm{S}_{i}$, and its Fourier mode $\mathbf{M}_{k}=L^{-2}\left|\sum_{i}\bm{S}_{i}e^{i\bm{k}\cdot\bm{r}_{i}}\right|$, where $\bm{k}=\frac{2\pi}{L}\hat{\bm{x}}$, the smallest wave vector of the reciprocal lattice in $\bm{x}$-direction.
@@
    $Q_{m}=\frac{\langle M^{2}\rangle^{2}}{\langle M^{4}\rangle}.$ (9)
@@
    The scaled covariance between $\varepsilon$ and $M^{2}$, defined as
    $K$ $=-\frac{L^{2}}{\langle M^{2}\rangle}\left(\langle\varepsilon M^{2}\rangle-\langle\varepsilon\rangle\langle M^{2}\rangle\right).$ (10)
  5. (v)
    The specific-heat-like quantity, defined as
    $C_{\text{NN}}=\beta^{2}L^{2}\left(\langle\varepsilon^{2}\rangle-\langle\varepsilon\rangle^{2}\right).$ (11)
    where $\beta=1/T$ is the inverse temperature.
Note that the second-moment correlation length $\xi_{\text{2nd}}$ is different from the commonly referred exponential correlation length $\xi_{\text{exp}}$. The former is defined based on the Fourier spectrum of the correlation function [41, 43], while the latter is obtained by fitting the exponentially decaying correlation function. Both correlation lengths are well-defined in the disordered phase and become asymptotically equivalent in the thermodynamic limit [44]. In practice, the second-moment correlation length $\xi_{\text{2nd}}$ can be easily measured in MC simulations, whereas determining $\xi_{\text{exp}}$ can be challenging because corrections simultaneously arising from both finite distance and finite system sizes and their mixing. In addition, near and at the critical point, $\xi_{\rm exp}$ becomes ill-defined since the correlation function exhibits a power-law scaling. In the long-range ordered phase, it is extremely challenging to extract $\xi_{\rm exp}$ since the correlation function saturates to a non-zero constant as the distance becomes infinitely large. Throughout this work, unless specified otherwise, the correlation length $\xi$ refers to the second-moment correlation length $\xi_{\text{2nd}}$, which is well defined in all disordered, critical, and long-range ordered phases.
The dimensionless ratio $\xi/L$ is an effective tool for identifying the critical points and the phase transition properties [45, 46, 42]. In the disordered phase, the correlation length $\xi$ is finite, and the ratio $\xi/L$ vanishes as $L$ increases. At criticality, $\xi$ is proportional to the linear system size $L$ from the algebraic decaying behavior of the correlation function, and the ratio $\xi/L$ takes a universal value in the thermodynamic limit. In the ordered phase, $\xi/L$ diverges since $M^{2}$ remains finite while $M_{k}^{2}$ vanishes with increasing $L$. As a consequence, for a continuous phase transition separating a long-range ordered from a disordered phase, the intersection of $\xi/L$ curves for different $L$ accurately pinpoints the critical point. In the case of a BKT transition, however, the ratio converges to a smooth and universal curve since the whole low-T QLRO phase is essentially critical with an algebraically decaying correlation function.
Throughout the simulation, the error bars are estimated using standard error analysis techniques: for simple observables, such as $\varepsilon,M^{2}$, and $M_{k}^{2}$, the error bars are calculated using the standard binning method; for composite observables obtained by the arithmetic combinations of simple observables, such as the 2nd-moment correlation length $\xi$ and Binder ratio $Q$, the error bars are estimated using the standard jackknife method.
@@ II.4 Finite-size Scaling Analysis
Finite-size scaling (FSS) analysis is an essential technique in numerical investigation of critical phenomena. In the language of the renormalization group (RG) theory, the critical behavior of systems belonging to the same universality class is governed by a common fixed point. Near critical points, the free energy, and thus other physical observables, can be expressed as universal functions of scaling fields. FSS extends these universal functions to finite systems, enabling precise determination of critical points and accurate extraction of critical exponents. Near the critical point, the free-energy density $f(t,h,u,L^{-1})$ is a function of the thermal scaling field $t$, the magnetic scaling field $h$, an irrelevant field $u$, and the finite-size field $L^{-1}$. Then, the FSS behavior of free-energy density is given by,
@@
where $y_{t}=1/\nu$ is the thermal renormalization exponent ($\nu$ is the correlation-length exponent); $y_{h}$ is the magnetic renormalization exponent; $y_{i}<0$ is the leading irrelevant scaling exponent; and $g$ is the analytic part of free-energy density. The scaling fields are non-universal functions of microscopic interactions and external physical parameters. To the lowest order, $t$ is proportional to $T-T_{c}$, i.e., the distance between temperature and critical temperature; $h$ is proportional to the physical external magnetic field $H$. The irrelevant field $u$ reflects the distance of criticality of the system and the fixed point.
The FSS behavior of various physical observables near $T_{c}$ can be derived from Eq. (12). Here, we present the scaling form of several key quantities: (a) magnetic susceptibility $\chi$, (b) scaled covariance $K$, and (c) dimensionless quantities, such as $Q_{m}$ and $\xi/L$.
(a) FSS of $\chi$. The magnetic susceptibility $\chi$ is calculated from the averaged squared magnetization $\langle M^{2}\rangle$, which is propositional to $\partial^{2}f/\partial h^{2}$ at $h=0$. Thus, near the critical point, the magnetic susceptibility scales as
@@
(b) FSS of $K$. The quantity $K$ is proportional to the correlation between the squared magnetization $M^{2}$ and the NN energy $\varepsilon$. The NN energy $\varepsilon$ shares a similar scaling behavior as the total energy $E$ near $T_{c}$ as $\sim L^{y_{t}-d}$ [39, 47]. Consequently, the correlation between $M^{2}$ and $\varepsilon$ is proportional to the temperature derivative of squared magnetization $\frac{\partial}{\partial t}\langle M^{2}\rangle$ [39]. The FSS form of $K$ at the critical point $t=0$ is then given by
@@
where $a_{i}$ are non-universal constants. The leading $L^{y_{t}}$ scaling behavior in Eq. (15) makes $K$ an excellent quantity to estimate the thermal scaling exponent $y_{t}=1/\nu$.
(c) FSS of dimensionless ratios. Dimensionless ratios, such as Binder ratio $Q_{m}$ and correlation length ratio $\xi/L$, have been widely used in FSS analysis to locate the critical point of phase transitions [45, 48, 46]. Near the critical point, the Binder ratio follows the scaling behavior
@@
III Overview of Results